Bailey is on the tenth frame of her recent bowling competition and she needs to pick up the last pin for a spare and the first place trophy. She rolls the 7.39 kg ball down the lane and it hits the 1.41 kg pin head on. The ball was moving at 7.9 m/s before the collision. The pin went flying forward at 1.7 m/s. Determine the post-collision speed of the ball.


1
Expert's answer
2022-04-15T16:40:19-0400

"m_1v_1=m_1u_1+m_2u_2,"

"u_1=\frac{m_1v_1-m_2u_2}{m_1}=7.6~\frac ms."